# Sequential estimation of disturbed aerodynamic flows from sparse measurements via a reduced latent space

### v1

@@ 4 Conclusion
A critical component in bridging sparse pressure observations with high-dimensional flow state estimation is the construction of a low-dimensional latent representation. However, as discussed in this study, the resulting latent space lacks clear physical interpretability, making it challenging to relate individual latent components to specific flow structures or phenomena. Future work may focus on enhancing interpretability—either through targeted regularization strategies or by modifying the autoencoder architecture—to promote more physically meaningful latent variables and improve transparency in the estimation process. It should be emphasized that this study explored an extreme scenario in which the forecast operator lacks long-time accuracy, demonstrating that even under such conditions, surface pressure measurements contain sufficient flow information to effectively correct the predicted states through filtering. However, the limitations imposed by unobservable latent directions—those that cannot be recovered from pressure data alone—can be mitigated by learning a more accurate forecast operator, which provides a stronger prior for the filtering process. As stated in our discussions before, incorporating history-aware models such as Recurrent Neural Networks (RNNs) significantly enhances the temporal prediction of latent states. While this approach departs from the classical Markovian assumption typically required in sequential filtering, filters like the EnKF can still be applied in weakly or non-Markovian settings. In such cases, although we would sacrifice the strict mathematical optimality of the Kalman update, the improved forecast model can compensate by providing richer dynamics and reducing reliance on measurement correction alone. Nonetheless, these memory-based approaches should be adopted with caution, particularly when interpretability and theoretical guarantees are important. Finally, this framework is broadly applicable to systems where a reduced-order state representation can be constructed.

### v2

@@ 4 Conclusion
A critical component in bridging sparse pressure observations with high-dimensional flow state estimation is the construction of a low-dimensional latent representation. However, as discussed in this study, the resulting latent space lacks its own physical interpretability without the use of the encoder to lift it back to the full flow space, making it challenging to relate individual latent components to specific flow structures or phenomena. Future work may focus on enhancing interpretability—either through targeted regularization strategies or by modifying the autoencoder architecture—to promote more physically meaningful latent variables and improve transparency in the estimation process. It should be emphasized that this study explored an extreme scenario in which the forecast operator lacks long-time accuracy, demonstrating that even under such conditions, surface pressure measurements contain sufficient flow information to effectively detect arbitrary gusts and correct the predicted states through filtering. However, the limitations imposed by weakly-observed latent directions—those that cannot be recovered from pressure data alone—can be mitigated by either learning a more accurate forecast operator, which provides a stronger prior for the filtering process, or by augmenting the measurement vector with other sensor data. This will be addressed in our future work. Finally, this framework is broadly applicable to systems where a reduced-order state representation can be constructed.
Appendix A Update step of the ensemble Kalman filter
For the update formula defined in Eq. (6), the Kalman gain from the standard (linear) Kalman filter is
$\boldsymbol{K}$ $=\boldsymbol{\Sigma}_{\boldsymbol{\xi}\boldsymbol{p}}\,\boldsymbol{\Sigma}_{\boldsymbol{p}\boldsymbol{p}}^{-1}$ (A.1)
$=\boldsymbol{\Sigma}_{f}\boldsymbol{H}^{\top}\left(\boldsymbol{H}\boldsymbol{\Sigma}_{f}\boldsymbol{H}^{\top}+\boldsymbol{\Sigma}_{\boldsymbol{\eta}}\right)^{-1},$ (A.2)
where $\boldsymbol{\Sigma}_{\boldsymbol{\xi}\boldsymbol{p}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{p}\boldsymbol{p}}$ represent the state-observation cross-covariance and observation covariance matrices, respectively, $\boldsymbol{\Sigma}_{f}$ is the prior covariance, $\boldsymbol{H}$ is the linear observation operator (or the tangent linear of a nonlinear observation operator, such as $\boldsymbol{h}_{\boldsymbol{W}}$ in the present study), and $\boldsymbol{\Sigma}_{\boldsymbol{\eta}}$ is the measurement-noise covariance defined in Section 2.3. Here, we aim to derive the ensemble approximation of these covariances.
We use the subscripts $f$ and $a$ to denote forecast (prior) and analysis (posterior) quantities, respectively. Given an ensemble of state $\{\boldsymbol{\xi}^{i}\}_{i=1}^{M}$ with $\boldsymbol{\xi}^{i}\in\mathbb{R}^{n}$, define the ensemble matrix $\boldsymbol{\Xi}=[\boldsymbol{\xi}^{1},\dots,\boldsymbol{\xi}^{M}]\in\mathbb{R}^{n\times M}$ and the ensemble mean
$\bar{\boldsymbol{\xi}}=\frac{1}{M}\sum_{i=1}^{M}\boldsymbol{\xi}^{i}.$ (A.3)
Let $\boldsymbol{\Xi}^{\prime}$ be the normalized anomaly matrix with columns
$\boldsymbol{\Xi}^{\prime}]^{i}=\frac{1}{\sqrt{M-1}}(\boldsymbol{\xi}^{i}-\bar{\boldsymbol{\xi}}).$ (A.4)
These definitions apply to both the forecast and analysis ensembles, yielding $\bar{\boldsymbol{\xi}}_{f}$, $\boldsymbol{\Xi}_{f}^{\prime}$, $\bar{\boldsymbol{\xi}}_{a}$, and $\boldsymbol{\Xi}_{a}^{\prime}$.
With the perturbed observation vector as defined in Eq. (5), where $\boldsymbol{\eta}^{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\boldsymbol{\eta}})$, let us define the innovation anomaly matrix with $i$th column given by
$\boldsymbol{P}^{\prime}_{f}]^{i}=\frac{\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi}_{f}^{i})+\boldsymbol{\eta}^{i}-\boldsymbol{h}_{\boldsymbol{W}}(\bar{\boldsymbol{\xi}}_{f})-\bar{\boldsymbol{\eta}}}{\sqrt{M-1}},$ (A.5)
with $\bar{\boldsymbol{\eta}}$ the sample mean of $\{\boldsymbol{\eta}^{i}\}$. It should be noted that this mean is itself a random number whose expected value is 0.
To connect the Kalman filter gain (A.1) to its ensemble form, we follow Asch et al. [1] and employ the standard linearization:
$\boldsymbol{H}(\boldsymbol{\xi}_{f}^{i}-\bar{\boldsymbol{\xi}}_{f})\simeq\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi}_{f}^{i})-\bar{\boldsymbol{p}}_{f}.$ (A.6)
Using this approximation, it is then easy to show that [1]:
$\boldsymbol{\Sigma}_{f}\boldsymbol{H}^{\top}\simeq\boldsymbol{\Xi}^{\prime}_{f}\boldsymbol{P}^{\prime\top}_{f}=\text{Cov}(\boldsymbol{\Xi}_{f},\boldsymbol{P}_{f}),$ (A.7)
and
$\left(\boldsymbol{H}\boldsymbol{\Sigma}_{f}\boldsymbol{H}^{\top}+\boldsymbol{\Sigma}_{\boldsymbol{\eta}}\right)\simeq\boldsymbol{P}^{\prime}_{f}\boldsymbol{P}^{\prime\top}_{f}=\text{Cov}(\boldsymbol{P}_{f},\boldsymbol{P}_{f}).$ (A.8)